### v1

@@ Abstract
Finding a few solutions for a given problem that are diverse, as opposed to finding a single best solution to solve the problem, has recently become a notable topic in theoretical computer science. It is important to mention that finding a single solution is computationally hard in most interesting cases. To overcome this issue, we analyze the problems under a finer scope of parameterized complexity, which allows for tractable algorithms in many cases. That in turn enables us to find a diverse set of solutions.
Recently, Baste, Fellows, Jaffke, Masařík, Oliveira, Philip, and Rosamond showed that under a standard structural parameterization by treewidth, one can find a set of diverse solutions for many problems with only a very small additional cost [Artificial Intelligence 2022]. In this paper, we investigate a much stronger graph parameter, the cliquewidth, which can additionally describe some dense graph classes. Broadly speaking, it describes graphs that can be recursively constructed by a few operations defined on graphs whose vertices are divided into a bounded number of groups while each such group behaves uniformly with respect to any operation.
We show that for any vertex problem, if we are given a dynamic program solving that problem on cliquewidth decomposition, we can modify it to produce a few solutions that are as diverse as possible with as little overhead as in the above-mentioned treewidth paper. As a consequence, we prove that a diverse version of any MSO1 expressible problem can be solved in FPT time parameterized by cliquewidth, the number of sought solutions, and the number of quantifiers in the formula. That was an important missing piece in the complexity landscape of structural graph parameters and logic. We prove our results allowing for a more general natural collection of diversity functions compared to only two mostly studied diversity functions previously. That might be of independent interest as a larger pool of different diversity functions can highlight various aspects of different solutions to a problem.
Keywords: parameterized complexity, combinatorial optimization, diverse solutions, cliquewidth, dynamic programming.
@@ 1 Introduction
In a standard graph problem, one typically asks for a solution of the smallest size. The first difficulty that typically needs to be overcome is that a large proportion of important problems is hard to solve (for some notion of hardness). There is a vast branch of research dedicated to overcoming such a difficulty, like parameterized complexity, approximation algorithms, and heuristics, among others. Second, maybe a more pressing issue is that even once we are given a single optimal solution for the problem, we might not be able to learn as much about our original tasks as we had hoped. Why it might be so? Mathematically formulated problems are typically just an abstraction of real-world problems, where there might be important other factors that were perhaps just too complex to formalize or even not known in the first place. Such constraints might be visible only once an actual solution is presented. While the first difficulty stands, the second difficulty might be overcome or at least diminished by the following idea. Instead of searching for just one solution, we will search for various solutions that are as different as possible (for some notion of what different means, to be specified later). Moreover, the idea of looking for a diverse set of solutions may even further allow us to relax the original mathematical formulation of a problem which in turn may allow for more efficient algorithms combating the first difficulty mentioned. There are many examples of advantages producing numerous different solutions in the literature; see the famous architect problem [Gal89] as one of the earliest examples. Since then there have been countless applications of this principle in various branches of computer science: constraint programming [HOW07, HHOW05, PT15], mixed integer optimization [AMT24, DW09, GLW00], genetic algorithms [DGNN22, GBPS18, WO03], social choice [BN21, AFL+21], and discrete geometry [Kv22].
It was not until 2019 that this concept was brought to the attention from the perspective of theoretical graph algorithms and parameterized complexity toolbox [BFJ+22]. They showed that a dynamic program (DP) on a tree decomposition can be converted to a DP finding diverse solutions. Since then this concept under the parameterized complexity examination started to play a more prominent role and many follow-up papers have been published [BJM+19, HKKO21, dLMS23, FGP+23, HKK+23, AFdW23, DGNN23, FGJ+24].
@@
Having defined the basic measures, we can state the problem formally. In the following, we use $Div$ to denote an arbitrary diversity measure that takes $r$ subsets of vertices and outputs an integer, for example, $DivSum$ or $DivMin$ functions defined above. Note that in this work, we are concerned only with vertex problems, but the template can be easily modified to take into account any list of subsets. Formaly, a vertex problem is a set of pairs $(G,S)$, where $G$ is a graph and $S$ is a subset of its vertices, that satisfies some property. For example, we can define a vertex problem $\mathcal{P}$ as the set of pairs $(G,S)$ such that $|S|\leq k$ and $S$ is a vertex cover of $G$.
@@
We say that graph $G$ has cliquewidth $\omega$ if there is a rooted tree that describes the construction of $G$ from single vertices via the following operations that are using at most $\omega$ colors at any given time during the construction. The operations are: disjoint union of two graphs, recolor all vertices of one color to another one, and add all edges with endpoints of specified colors. Check Section 2.2 for a formal definition.
The cliquewidth has played an influential role since its definition in 1993 [CER93]. This has only increased after a seminal result of Courcelle, Makowsky, and Rotics proving that any MSO1 property can be checked in linear time on graphs of bounded cliquewidth [CMR00]. This result essentially means that many difficult vertex problems on graphs are solvable in linear time provided that the cliquewidth of that graph is small. The most prominent tool to solve problems on graphs of bounded cliquewidth is dynamic programming without any doubts. Besides the model-checking theorem, there have been countless examples of using dynamic programming to solve some particular problems. For example, feedback vertex set [BXSTV13], triangle detection & girth [CDP19], Alliances [KO17], Hamiltonian cycle [BKK19], Steiner tree [BK24b], connected odd cycle transversal [BK24a], and other connectivity problems [BK19, HK23].
We would like to point out that for our purposes, there is very little difference between the case when the cliquewidth decomposition is given on the input and the case when it first needs to be computed. Indeed, despite being NP-hard to be computed optimally [FRRS09] there is a $(2^{2k+1}-1)$-approximation algorithm running in quadratic time [FK22]. Therefore, from now on we will simply assume that the decomposition is given on the input, so we do not need to argue this in many places. Moreover, it allows us to state our results independently from potential improvements of the above-mentioned approximation algorithm.
@@ 1.1 Our Results
More formally, we define what means for dynamic programming to be monotone; see Definition 4.5 for a formal definition. Despite its inherent technical nature, the idea behind this definition is very natural. Let us explain the intuition here. Take a dynamic programming table entry $\varepsilon$ within some inner node of the DP. Such $\varepsilon$ could originate as many different possible combinations of other entries evaluated earlier (in subtrees of the given node) in the DP. We call it the derivation tree of $\varepsilon$. To each such derivation tree, we assign a set of vertices, called partial solution, based only on the entries in the leaves of the decomposition that combine to $\varepsilon$. Moreover, for the leaf node, we only allow for two entries: one where the vertex corresponding to a leaf node is in the partial solution and the other when it is not. We call the described property monotonicity, as the set of vertices assigned to the derivation tree is the union of the sets assigned to its subtrees.
With the intuition on how monotone DPs look like, we can formulate our main theorem. In the following, we use $Div$ to denote an arbitrary diversity measure from a collection of Venn measures that will be defined later; see Section 2.3 for a formal definition. At this point, we can keep in mind that $DivSum$ is one such measure, but $DivMin$ is not.
@@
Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{r}$ be vertex problems and given monotone dynamic programs solving them with the slowest one running in time $t_{\mathcal{D}}$ on graph $G$ with cliquewidth decomposition $\mathcal{D}$. Let $Div$ be a Venn diversity function. The problem $Div^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$ on a graph $G$ with given cliquewidth decomposition $\mathcal{D}$ can be solved in time $\mathcal{O}(t_{\mathcal{D}}^{r})$.
It is important to note that the running time of Theorem 1.2 does not depend on the actual value of the $Div$ function. Our proof provides the same statement also for the $DivMin$ function, but with a worse running time dependent on the target value $d$ of the $DivMin$ function. In particular, the running time will be $\mathcal{O}(t_{\mathcal{D}}^{r}d^{r^{2}})$ in that case.
@@
Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{r}$ be vertex problems and given monotone dynamic programs solving them with the slowest one running in time $t_{\mathcal{D}}$ on graph $G$ with cliquewidth decomposition $\mathcal{D}$. The problem $DivMin^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$ on a graph $G$ with given cliquewidth decomposition $\mathcal{D}$ can be solved in time $\mathcal{O}(t_{\mathcal{D}}^{r}d^{r^{2}})$.
First, we would like to note that factor $d^{a}$ (where typically $a=2$) stated in the running time of [BFJ+22, Theorem 10] is not necessary. The reason for this is that during the DP composition, we need to remember for each entry only the best value of diversity and not all of them similarly as in our proof of Theorem 1.2. Therefore, a reformulated statement of [BFJ+22, Theorem 10] with the above-mentioned improvement follows which also serves as a comparison that our algorithm qualitatively matches the running time of the previous work.
@@
Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{r}$ be vertex problems and given dynamic program core (as defined in [BFJ+22, Definition 6]) solving them with the slowest one running in time $t_{\mathcal{T}}$ on graph $G$ with tree decomposition $\mathcal{T}$. The problem $DivSum^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$ on a graph $G$ with given treewidth decomposition $\mathcal{T}$ can be solved in time $\mathcal{O}(t_{\mathcal{T}}^{r})$.
We would like to argue now that a vast majority of dynamic programs solving vertex problems using the cliquewidth decomposition are either monotone or could be easily transformed to be monotone. To our best knowledge, we have not found any examples of non-monotone DPs in the literature. To support the intuition, we designed a non-monotone DP for the vertex cover problem that we explain in detail in Section 4.4. However, this algorithm can be easily turned into a monotone one, which we present in Section 3.
In order to support our claim, we showed that any problem that can be expressed in MSO1 logic can be indeed solved in the diverse version within our framework. In other words, we show that MSO1 vertex problems can be solved by a monotone dynamic program to which we apply Theorem 1.2. As the original result [CMR00] gives the DP only implicitly we decided to design a new DP algorithm and show that it is indeed monotone. We based our approach on lecture notes [Fia24, Section 5.4] which shows how to use evaluation trees to obtain a similar result on treewidth decomposition.
@@
1.1.1 New Venn Diversity Measures.
In our work, we propose a whole collection of diversity measures. We believe they will be of an independent interest. Despite being very natural, we have not found them studied in this context before up to our best knowledge. In the previous theoretical works, research was only focused on the two measures already defined: $DivMin$ and $DivSum$. However, despite them being first natural examples there is nothing that special about them. One can find examples where those measures do not behave exactly as expected. In the case of the $DivSum$ measure, a profound example is taking many copies of two disjoint sets which results in a relatively high diversity value, but intuitively such a tuple of sets is not diverse. In particular, in such a case such a diverse solution does not display a potentially rich structure of various different solutions as we might have hoped. This was already observed and discussed in [BJM+19, Section 5, Figure 3]. Such problems are not limited to $DivSum$. It is easy to imagine a similar situation with the $DivMin$ measure. Indeed, once the number of solutions we seek is larger than the number of different solutions that differ from each other on at least $d$ vertices, we only receive information that no such diverse set exists. At first, this might look like having an easy fix simply by decreasing our diversity demand $d$. However, there are instances where doing so leads only to outputs that have very little difference. In particular, think of the solution being everywhere the same except for a few ($d$) vertices, while the whole structure may contain many quite distinct solutions. Similarly to the examples above, we think that for each potential diversity measure, some problematic instances can be created so that when the measure is used the diverse collection does not reflect at all how diverse the actual solutions might be.
Therefore, we believe that our approach which gives access to the whole collection of diversity measures much better supports the nature of the diverse problems. Indeed, it seems in line with the diversity problem where one seeks a collection of solutions rather than a single one of them. In the same spirit, we advocate to try different different diversity measures and then compare their results. However, we can do even more when having access to more measures. We can choose the measure based on the instance and based on what results we have obtained so far. If we obtain one of the degenerate examples highlighted above or simply if we do not like the solution given by one diversity function, we can adjust the diversity function to penalize such an outcome.
We will shortly give an example of how one can do it, but first, we define Venn diversity measures formally and give an example of one such measure. We start by defining the membership vector which is basically a characteristic vector of a vertex $v$ being an element of set $S_{i}$ in the list of sets $S_{1},\ldots,S_{r}$.
@@
For $v\in V$ and a list $S_{1},\ldots,S_{r}$ of subsets of $V$ a membership vector of $v$ in $S_{1},\ldots,S_{r}$ is a vector $m(S_{1},\ldots,S_{r},v)\in\{0,1\}^{r}$ such that for each $i\in[r]$
@@
Then we define an influence of a vertex provided with function $f$ where the value of $f$ is determined by the membership vector value.
@@
Observe that in this definition the influence depends only on the presence of $v$ in each of $S_{1},\ldots,S_{r}$. Finally, we define the Venn function.
@@
$Div_{f}(S_{1},\ldots,S_{r})\coloneqq\sum_{v\in V}I_{f}(v).$
Whenever $f$ is fixed, we will write $I(S_{1},\ldots,S_{r},v)$ instead of $I_{f}(S_{1},\ldots,S_{r},v)$ and $Div(S_{1},\ldots,S_{r})$ instead of $Div_{f}(S_{1},\ldots,S_{r})$, and we say Venn diversity instead of Venn $f$-diversity.
We now give an example of a Venn $f$-diversity measure other than $DivSum$ (a formal argument that $DivSum$ is a Venn $f$-diversity measure is postponed to Section 2.3 ).
Example 1.9.
@@
Here the influence of $v$ on $S_{1},\ldots,S_{r}$ depends on the intersection of how many sets from $S_{1},\ldots,S_{r}$ $v$ belongs to.
Returning to our example with only two distinct solutions copied multiple times achieving the best $DivSum$ value, we suggest that in that case one might try to use the $Div^{*}$ measure defined in Example 1.9. Contrary to the sum of pairwise hamming distances, this measure penalizes taking many copies of two disjoint sets. See Figure 1 for an example of such a behavior.
@@
Figure 1: Consider the dominating set problem on a path on five vertices. If the number of required solutions is four, then a solution maximizing $DivSum$ will consist of twice A and twice B, having $DivSum$ equal to 20. However, for the $Div*$ function, the value for this set of solutions is equal to 60, which is less than for a more diverse set of solutions consisting of all A, B, C, and D, with the $Div*$ being equal to 63.
@@
After short preliminaries (Section 2) we present an exposition of our main proof and tools on an example of the vertex cover problem (Section 3). Then we give the proof of our main theorem (Theorem 1.2) in Section 4. We present an example of a non-monotone dynamic program in Section 4.4. Then we include a section showing monotone DP for MSO1 problems Section 5. That provides a proof of Corollary 1.5. We conclude with a section containing a discussion about open problems and further extension Section 6.

### v2

@@ Abstract
Finding a few solutions for a given problem that are diverse, as opposed to finding a single best solution to solve the problem, has recently become a notable topic in theoretical computer science. It is important to mention that finding a single solution is computationally hard in most interesting cases. To overcome this issue, we analyze the problems under a finer scope of parameterized complexity, which often allows for tractable algorithms. That, in turn, enables us to find a diverse set of solutions.
Recently, Baste, Fellows, Jaffke, Masařík, Oliveira, Philip, and Rosamond showed that under a standard structural parameterization by treewidth, one can find a set of diverse solutions for many problems with only a very small additional cost [Artificial Intelligence 2022]. In this paper, we investigate a much stronger graph parameter, the cliquewidth, which can additionally describe some dense graph classes. Broadly speaking, it describes graphs that can be recursively constructed by a few operations defined on graphs whose vertices are divided into a bounded number of groups, while each such group behaves uniformly with respect to any operation.
We show that for any vertex problem, if we are given a dynamic program solving that problem on cliquewidth decomposition, we can modify it to produce a few solutions that are as diverse as possible with as little overhead as in the above-mentioned treewidth paper. As a consequence, we prove that a diverse version of any MSO1 expressible problem can be solved in linear FPT time parameterized by the cliquewidth, the number of sought solutions, and the number of quantifiers in the formula, which was a natural missing piece in the complexity landscape of structural graph parameters and logic for the diverse problems. We prove our results, allowing for a more general natural collection of diversity functions compared to only two mostly studied diversity functions previously. That might be of independent interest as a larger pool of different diversity functions can highlight various aspects of different solutions to a problem.
Keywords: Parameterized complexity, combinatorial optimization, diverse solutions, diversity measures, cliquewidth, dynamic programming, MSO logic.
@@ 1 Introduction
In a standard graph problem, one typically asks for a solution of the smallest size. The first difficulty that typically needs to be overcome is that a large proportion of important problems is hard to solve (for some notion of hardness). There is a vast branch of research dedicated to overcoming such a difficulty, like parameterized complexity, approximation algorithms, and heuristics, among others. Second, maybe a more pressing issue is that even once we are given a single optimal solution for the problem, we might not be able to learn as much about our original tasks as we had hoped. Why might it be so? Mathematically formulated problems are typically just an abstraction of real-world problems, where there might be other important factors that were perhaps just too complex to formalize or even not known in the first place. Such constraints might be visible only once an actual solution is presented. While the first difficulty stands, the second difficulty might be overcome or at least diminished by the following idea. Instead of searching for just one solution, we will search for various solutions that are as different as possible (for some notion of what different means, to be specified later). Moreover, the idea of looking for a diverse set of solutions may even further allow us to relax the original mathematical formulation of a problem, which in turn may allow for more efficient algorithms combating the first-mentioned difficulty. There are many examples of advantages producing numerous different solutions in the literature; see the famous architect problem [Gal89] as one of the earliest examples. Since then, there have been countless applications of this principle in various branches of computer science: constraint programming [HOW07, HHOW05, PT15, IGST20], mixed integer optimization [AMT24, DW09, GLW00], evolutionary computation [DGNN22, GBPS18, WO03], social choice [BN21, AFL+21], and discrete geometry [Kv22] among many others.
It was not until 2019 that this concept was brought to the attention from the perspective of theoretical graph algorithms and parameterized complexity toolbox [BFJ+22]. They showed that a dynamic program (DP) on a tree decomposition can be converted to a DP finding diverse solutions. Since then, this concept under the parameterized complexity examination started to play a more prominent role, and many follow-up papers have been published extending our knowledge into various directions: diverse hitting sets [BJM+19], diverse polynomial-time solvable problems (spanning trees, paths) [HKKO21], diverse minimum $st$-cuts [dLS23, dLS25, IMMS25], diversity in matroids [FGP+24] approximation algorithms framework for diverse solutions [HKK+23], diverse synchronization of words [AFOW23], Submodular function approximations and experiments [DGNN23], diverse SAT [MMR24], and diverse pairs of matchings [FGJ+24].
@@
Having defined the basic measures, we can state the problem formally. In the following, we use $Div$ to denote an arbitrary diversity measure that takes $r$ subsets of vertices and outputs an integer, for example, $DivSum$ or $DivMin$ functions defined above. Note that in this work, we are concerned only with vertex problems, but this template can be easily modified to take into account any list of subsets. Formally, a vertex problem is a set of pairs $(G,S)$, where $G$ is a graph and $S$ is a subset of its vertices that satisfies some property. For example, we can define k-Vertex Cover as the set of pairs $(G,S)$ such that $|S|\leq k$ and $S$ is a vertex cover of $G$.
@@
In our paper, we focus on FPT algorithms for diverse vertex problems, parameterized by the cliquewidth of the graph.
A graph $G$ has cliquewidth $\omega$ if there is a rooted tree that describes the construction of $G$ from single vertices via the following operations that use at most $\omega$ colors at any time during the construction. The operations are: disjoint union of two graphs, recolor all vertices of one color to another one, and add all edges with endpoints of specified colors. Inspect Section˜2.2 for a formal exposition.
The cliquewidth has played an influential role since its definition in 1993 [CER93]. This has only increased after a seminal result of Courcelle, Makowsky, and Rotics proving that any MSO1 property can be checked in linear time on graphs of bounded cliquewidth [CMR00]. This result essentially means that many difficult vertex problems on graphs are solvable in linear time, provided that the cliquewidth of that graph is small. The most prominent tool to solve problems on graphs of bounded cliquewidth is dynamic programming without any doubt. Besides the model-checking theorem, there have been countless examples of using dynamic programming to solve some particular problems, either because those problems are not expressible in MSO1 or because they provide a much faster algorithm. For example, feedback vertex set [BSTV13], triangle detection & girth [CDP19], Alliances [KO17], Hamiltonian cycle [BKK19], Steiner tree [BK24], connected odd cycle transversal [BK25], and other connectivity problems [BK19, HK23].
We would like to point out that for our purposes, there is very little difference between the case when the cliquewidth decomposition is given on the input and the case when it first needs to be computed. Indeed, despite being NP-hard to be computed optimally [FRRS09], there is a $(2^{2\omega+1}-1)$-approximation algorithm running in quadratic time, where $\omega$ is the cliquewidth of the graph on the input [FK24]. Therefore, from now on, we will simply assume that the decomposition is given on the input, so we do not need to argue this in many places. Moreover, it allows us to state our results independently of potential improvements of the above-mentioned approximation algorithm.
@@ 1.1 Our Results
More formally, we define what it means for dynamic programming to be monotone; see Definition˜3.5 for a formal definition. Despite its inherent technical nature, the idea behind this definition is very natural. Let us explain the intuition here. Take a dynamic programming table entry $\varepsilon$ within some inner node of the DP. Such $\varepsilon$ could originate from many different possible combinations of other entries evaluated earlier (in subtrees of the given node) in the DP. We call it the derivation tree of $\varepsilon$. To each such derivation tree, we assign a set of vertices, called partial solution, based only on the entries in the leaves of the decomposition that combine to $\varepsilon$. Moreover, for the leaf node, we only allow for two entries: one where the vertex corresponding to a leaf node is in the partial solution and the other when it is not. We call the described property monotonicity, as the set of vertices assigned to the derivation tree is the union of the sets assigned to its subtrees.
With the intuition on what monotone DPs look like, we formulate our main theorem. In the following, we use $Div$ to denote an arbitrary diversity measure from a collection of Venn measures that will be defined later; see Definition˜1.9 for a formal definition. At this point, we keep in mind that $DivSum$ is one such measure, but $DivMin$ is not.
@@
Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{r}$ be vertex problems and given monotone dynamic programs solving them with the slowest running on a single node of the decomposition in time $t_{\mathcal{D}}$ for graph $G$ with cliquewidth decomposition $\mathcal{D}$. Let $Div$ be a Venn diversity function. The problem $Div^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$ on a graph $G$ with given cliquewidth decomposition $\mathcal{D}$ can be solved in time $\mathcal{O}(|V(\mathcal{D})|t_{\mathcal{D}}^{r})$.
It is important to note that the running time of Theorem˜1.2 does not depend on the actual value of the $Div$ function. Our proof also provides the same statement for the $DivMin$ function, but with a worse running time dependent on the target value $d$ of the $DivMin$ function. In particular, the running time will be $\mathcal{O}(|V(\mathcal{D})|t_{\mathcal{D}}^{r}d^{r^{2}})$ in that case.
@@
Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{r}$ be vertex problems and given monotone dynamic programs solving them with the slowest running on a single node of the decomposition in time $t_{\mathcal{D}}$ for graph $G$ with cliquewidth decomposition $\mathcal{D}$. The problem $DivMin^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$ on a graph $G$ with given cliquewidth decomposition $\mathcal{D}$ can be solved in time $\mathcal{O}(|V(\mathcal{D})|t_{\mathcal{D}}^{r}d^{r^{2}})$.
We would like to compare the running time of Theorem˜1.2 with an analogous theorem for the treewidth [BFJ+22, Theorem 10]. We present it with a minor improvement, where we omit the factor $d^{a}$ (typically a=2{a=2}) stated in the referenced running time as it is not necessary once one aims to find a single diverse solution of the best possible value of $DivSum$. We discuss the change at the end of Section˜3.
@@
Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{r}$ be vertex problems and given dynamic program core (as defined in [BFJ+22, Definition 6]) solving them with the slowest running on a single node of the decomposition in time $t_{\mathcal{T}}$ for graph $G$ with tree decomposition $\mathcal{T}$. The problem $DivSum^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$ on a graph $G$ with given treewidth decomposition $\mathcal{T}$ can be solved in time $\mathcal{O}(|V(\mathcal{T})|t_{\mathcal{T}}^{r})$.
We claim that a vast majority of dynamic programs solving vertex problems using the cliquewidth decomposition are either monotone or could be easily transformed to be monotone. To the best of our knowledge, we have not found any examples of non-monotone DPs in the literature. To support the intuition, we designed a non-monotone DP for the vertex cover problem; see Section˜4.4. However, this algorithm can be easily turned into a monotone one, which we present in Section˜4.1. In order to support our claim, we showed that using our framework, one can solve the diverse version of any vertex problem that can be expressed in MSO1 logic, which is the fragment of second-order logic where the second-order quantification is limited to quantifying over sets of vertices. The MSO1 captures many well-studied problems, including: Vertex Cover, Independent Set, Dominating Set, and Feedback Vertex Set. We show that any problem $\mathcal{P}$ for which there exists an MSO1 formula $\Phi(S)$ such that $G\vDash\Phi(S)$ if and only if $(G,S)\in\mathcal{P}$ can be solved by a monotone dynamic program to which we apply Theorem˜1.2. As the original result [CMR00] gives the DP only implicitly, we decided to design a new DP algorithm and show that it is indeed monotone. We based our approach on lecture notes [Fia24, Section 5.4] [^1] This approach originates in an unpublished work of Hliněný, Král’, and Obdržálek., which show how to use evaluation trees to obtain a similar result for treewidth decomposition.
@@
Corollary˜1.5 is interesting on its own as it provides a new result for graph vertex problems expressible in certain logic. We can compare it with the original treewidth result [BFJ+22], which encompasses problems that can be defined in MSO2 logic and then provides the diverse variant in the running time comparable to Corollary˜1.5. That is a stronger logic, but a much weaker graph class. Note that model checking of MSO2 problems on cliques is not even in XP unless $\text{E}=\text{NE}$ [CMR00]. Later, Hanaka, Kobayashi, Kurita, and Otachi [HKKO21, Theorem 13] proved that problems that can be defined in FO logic on a nowhere dense graph class can have their diverse variant solved in FPT time. However, this time the result relies heavily on the model-checking machinery of Grohe, Kreutzer, and Siebertz [GKS17] and the claimed running time is much worse than linear. Again, this result is incomparable with Corollary˜1.5 as it acts on a larger class of graphs, but only on problems defined in a weaker logic. In 2023, Bergougnoux, Dreier, Jaffke [BDJ23] worked on an A&C DN logic (distance neighborhood logic with acyclicity and connectivity constraints). Informally speaking, DN logic allows for size measurement of neighborhood terms as well as inclusion-wise comparison of neighborhoods within the existential fragment of MSO1 logic, while A&C further add two more constraints which allow us to express acyclicity and connectivity of the neighborhood term. See [BDJ23, Section 3] for a formal definition. They showed that if a vertex problem can be expressed in A&C DN logic, then one can express $DivMin$ or $DivSum$ variant of the problem in that logic as well [BDJ23, Observation 1.1]. Then they present a cubic FPT algorithm solving A&C DN encoded problems on graphs of bounded cliquewidth parameterized by the size of the formula and the cliquewidth. Additionally, they present an XP algorithm for the same problem on graphs of bounded mimwidth [^2] Informally, the mimwidth of a graph minimizes the size of each induced matching appearing in some particular cuts of the graph which are specified by a recursive decomposition of its vertex set; see [Vat12] for a formal definition. which is a much more general structural graph parameter. Indeed, as this covers only an existential fragment of MSO1, it is incomparable with Corollary˜1.5. Let us stress that all the related work is stated only for $DivSum$ or possibly for $DivMin$ measures.
New Venn Diversity Measures.
In our work, we propose a whole collection of diversity measures. We believe they will be of independent interest. Despite being very natural, to the best of our knowledge, they have not been studied in this context before. In the previous theoretical works, research was only focused on the two measures already defined: $DivMin$ and $DivSum$. However, despite them being the first natural examples, there is nothing so special about them. One can find examples where those measures do not behave exactly as expected. In the case of the $DivSum$ measure, a profound example is taking many copies of two disjoint sets, which results in a relatively high diversity value. Still, intuitively, such a tuple of sets is not diverse. In particular, it does not display a potentially rich structure of various different solutions as we might have hoped. This was already observed and discussed in [BJM+19, Section 5, Figure 3]; see also an example illustrated later in Figure˜3. Such problems are not limited to $DivSum$. See Example˜1.6 (Figure˜1), where the optimal $DivMin$ value is $2$ when seeking six solutions, while there is another tuple of six solutions with the same $DivMin$ value, which seems intuitively much more diverse. Indeed, a straightforward generalization of Example˜1.6 gives that $DivSum$ of the first $r$-tuple of solutions equals $2\binom{r}{2}$ and the second $r$-tuple has $DivSum$ arbitrarily high (depending only on $s$, as $r$ is fixed) while they both have the same $DivMin=2$. Similarly to the examples above, we believe that for each potential diversity measure, some problematic instances can be created so that when the specific measure is used, the most diverse collection does not reflect at all how diverse the whole space of solutions might be.
@@
Figure 1: Illustration of Example˜1.6 drawn as a bipartite incidence graph. Each feasible solution corresponds to an edge $(b_{j},A_{i})$ representing the set $A_{i}\cup\{b_{j}\}$.
Example 1.6.
We fix the sought number of solutions $r=6$. We partition the vertices into the following sets $B,A_{1},A_{2},A_{3},A_{4},A_{5}$, each of the same size $s\geq 6$. Consider a problem where only the following sets are possible solutions: any single vertex in $B$ together with the whole set $A_{i}$, $i\in[5]$. Then a possible set of six solutions with the maximum value of DivMinDivMin (=2=2) consists of sets $A_{1},b_{j}$, where $j\in[6]$. Such a solution has $DivSum$ equals $30$. However, there is a set of six solutions with intuitively bigger diversity (and also $DivSum=28s+30$) while its $DivMin$ remains at $2$. For example, take $A_{i},b_{i}$ for $i\in[5]$ and $A_{1},b_{6}$. See Figure˜1 depicting this example.
Therefore, we are convinced that our approach, which gives access to the whole collection of diversity measures, much better supports the nature of the diverse problems. Indeed, it seems in line with the diversity paradigm, where one seeks a collection of solutions rather than a single one. In the same spirit, we advocate trying different diversity measures and then comparing their results. However, we can be more subtle as we have access to a large collection of measures. We suggest that one might pick the measures dynamically. This means we can choose the measures not only based on the instance but also based on the diverse collection(s) we have obtained so far. If we obtain one of the degenerate examples, or simply, whenever we do not like the solution given by one diversity function, we can adjust the diversity function to penalize such an outcome and rerun the algorithm.
We will shortly give an example of a measure that penalizes an unwanted outcome returned using the $DivSum$ measure, but first, we define Venn diversity measures formally. We start by defining the membership vector, which is basically a characteristic vector of a vertex $v$ being an element of set $S_{i}$ in the list of sets $S_{1},\ldots,S_{r}$.
@@
For $v\in V$ and a list $S_{1},\ldots,S_{r}$ of subsets of $V$ a membership vector of $v$ in $S_{1},\ldots,S_{r}$ is a vector $m(S_{1},\ldots,S_{r},v)\in\{0,1\}^{r}$ such that for each $i\in[r]$,
@@
Then we define the influence of a vertex provided with a function $f$ where the value of $f$ is determined by the membership vector value.
@@
Observe that in this definition, the influence depends only on the presence of $v$ in each of $S_{1},\ldots,S_{r}$. Finally, we define the Venn function; see Figure˜2.
@@
$Div_{f}(S_{1},\ldots,S_{r})\coloneqq\sum_{v\in V}I_{f}(S_{1},\ldots,S_{r},v).$
AABBCCDD3333444444333340
Figure 2: A Venn diagram for four solutions: $A,B,C,D$. The influence of a vertex corresponding to the $DivSum$ measure is marked. It is the number of solutions it belongs to multiplied by the number of solutions it does not belong to. With Venn diversity measures, one could assign an arbitrary influence value to each part of this diagram.
Whenever $f$ is fixed, we will write $I(S_{1},\ldots,S_{r},v)$ instead of $I_{f}(S_{1},\ldots,S_{r},v)$ and $Div(S_{1},\ldots,S_{r})$ instead of $Div_{f}(S_{1},\ldots,S_{r})$, and we say Venn diversity instead of Venn $f$-diversity. We now give an example of a Venn $f$-diversity measure other than $DivSum$.
Example 1.10 ($Div^{*}$ measure).
@@
Returning to our example with only two distinct solutions copied multiple times, achieving the best $DivSum$ value, we suggest that in that case, one might try to use the $Div^{*}$ measure defined in Example˜1.10. Contrary to the sum of pairwise hamming distances, this measure penalizes taking many copies of two disjoint sets; see Figure˜3 for an example of a small problem instance already admiting such a behavior.
@@
Figure 3: Consider the dominating set problem on a path of five vertices. If the number of required solutions is four, then a solution maximizing DivSumDivSum will consist of twice A and twice B, having DivSumDivSum equal to 20. However, for the $Div^{*}$ function, the value for this set of solutions is equal to 60, which is less than for an arguably more diverse set of solutions consisting of all A, B, C, and D, with the $Div^{*}$ being equal to 63.
@@
After short preliminaries (Section˜2), we present the proof of our main theorem (Theorem˜1.2) in Section˜3. Then we provide an exposition of our main proof and tools on an example of the vertex cover problem (Section˜4). We also give a simple example of a non-monotone dynamic program in Section˜4.4. Then we include a section showing monotone DP for MSO1 problems Section˜5. That section provides a proof of Corollary˜1.5. We conclude with a discussion about open problems and further extensions in Section˜6.